# Relative Rényi Entropy Under Local Quenches in 2D CFTs

### v1

@@ 2 Setup in 2D CFTs
[…]
@@
3 Relative Rényi entropy for different states
In RCFTs, it is known that the excess of Rényi entropy for primary and descendant operators [27, 51, 28] and the pseudo-Rényi entropy for primary and descendant operators [55, 30] both saturate to a constant value. This constant is equal to the logarithm of the quantum dimension of the associated primary operator. Specifically, for two density matrices $\rho$ and $\sigma$ constructed from a primary operator $\mathcal{O}$ and a descendant operator $L_{-n}\mathcal{O}$, respectively, the entanglement or pseudo-entropies are identical when considering their time evolution. This result implies that primary and descendant operators are indistinguishable regarding entanglement and pseudo-entropy. Relative entropy, proposed as a measure of the distinguishability between two states, plays a central role in quantum information theory [36]. Therefore, this section will examine RRE for different states excited at the same position within the same conformal family in RCFTs to investigate how information-theoretic quantities of descendants might be discerned. For simplicity, we will choose an interval on the right of the inserted operator, noting that a conformal transformation allows us to obtain any desired insertion position.
Let us initially explore the most straightforward case that one density matrix is constructed from a vacuum state where we set $\mathcal{O}^{\prime}$ equals to the identity operator $\mathbbm{1}$ while the other is built from a primary operator $\mathcal{O}$, i.e.
@@ 3 Relative Rényi entropy for different states
The first term on the RHS of (9) is the increase of Rényi entropy of $\mathcal{O}$, and its evolution has been discussed exclusively in [27], which would equal to 00 at the early time ($t<\left|x\right|$) while equal to $\log d_{\mathcal{O}}$, where $d_{\mathcal{O}}$ is the quantum dimension of $\mathcal{O}$, at the late time ($t>\left|x\right|$). The second term on the RHS of (9), by utilizing the conformal map (5), can be reformulated as
$\text{tr}(\rho\sigma^{k-1})=\log\frac{\left|\frac{dw_{1}}{dz_{1}}\right|^{-2\Delta}\left|\frac{dw_{2}}{dz_{2}}\right|^{-2\Delta}\langle\mathcal{O}(z_{1},\bar{z}_{1})\mathcal{O}(z_{2},\bar{z}_{2})\mathbbm{1}\dots\mathbbm{1}\rangle_{\Sigma_{1}}}{\langle\mathcal{O}(w_{1},\bar{w}_{1})\mathcal{O}(w_{2},\bar{w}_{1})\rangle_{\Sigma_{1}}},$ (10)
It should be noted that Eq. (10) is supposed to have some contributions from the vacuum. However, since this part cancels out with the vacuum contribution in the entanglement entropy (4), (10) is just a convenient notation. At the early time, we have
@@
$\text{tr}(\rho\sigma^{k-1})=\log\frac{k^{-4\Delta}(x^{2}-t^{2})^{\frac{-2(k-1)\Delta}{k}}k^{4\Delta}(x^{2}-t^{2})^{\frac{2(k-1)\Delta}{k}}\left|2\epsilon\right|^{-4\Delta}}{\left|2\epsilon\right|^{-4\Delta}}=0.$ (12)
So, the RRE of the primary and identity operators is zero at the early time.
Next, we would discuss the late time behavior of (10). The numerator of the logarithm function can be evaluated as
@@
Therefore, at the late time limit, the $k^{\rm th}$ RRE of primary operator and identity operator saturates to
@@
In this subsection, we will discuss the time evolution of the RRE in more complex scenarios. The two density matrices are constructed from a primary operator and a descendant operator belonging to the same conformal family, i.e.
@@
Similar to (10), the notation of the form $\text{tr}(\rho\sigma^{k-1})$ with the definition of RRE (4) in this case is
$\text{tr}(\rho\sigma^{k-1})=\log\frac{\langle\mathcal{O}(w_{1},\bar{w}_{1})\mathcal{O}(w_{2},\bar{w}_{2})L_{-n}\mathcal{O}(w_{3},\bar{w}_{3})\dots L_{-n}\mathcal{O}(w_{2k},\bar{w}_{2k})\rangle_{\Sigma_{k}}}{\langle\mathcal{O}(w_{1},\bar{w}_{1})\mathcal{O}(w_{2},\bar{w}_{2})\rangle_{\Sigma_{1}}\langle L_{-n}\mathcal{O}(w_{1},\bar{w}_{1})L_{-n}\mathcal{O}(w_{2},\bar{w}_{2})\rangle_{\Sigma_{1}}^{k-1}}.$ (17)
@@
From (18) and (19) and replacing all the coordinates with (7), at the early time, we find the two-point functions on $\Sigma_{k}$ and on $\Sigma_{1}$ are the same, i.e.
@@
Therefore, combining (11), (17), (18) and (22), the $k^{\rm th}$ RRE of the primary operator and descendant operator at the early time is
@@
[…]
@@
$\sim$ $\langle\mathcal{O}(z_{1})L_{-n}\mathcal{O}(z_{4})\dots L_{-n}\mathcal{O}(z_{2j+1})L_{-n}\mathcal{O}(z_{2j+4})\dots L_{-n}\mathcal{O}(z_{2k-1})\mathcal{O}(z_{2})\rangle_{\Sigma_{1}}$
$\times$ $\langle\mathcal{O}(\bar{z}_{1})\mathcal{O}(\bar{z}_{2})\mathcal{O}(\bar{z}_{3})\dots\mathcal{O}(\bar{z}_{2k})\rangle_{\Sigma_{1}}+\dots$
@@
×$\times\langle\mathcal{O}(\bar{z}_{1})\mathcal{O}(\bar{z}_{2})\rangle_{\Sigma_{1}}\dots\langle\mathcal{O}(\bar{z}_{2k-3})\mathcal{O}(\bar{z}_{2k-2})\rangle_{\Sigma_{1}}\langle\mathcal{O}(\bar{z}_{2k-1})\mathcal{O}(\bar{z}_{2k})\rangle_{\Sigma_{1}}$ (25)
@@
As pointed out in [30], at the late time, the holomorphic part of the two-point function for two operators belonging to the same conformal family on $\Sigma_{k}$ and $\Sigma_{1}$ have the following relation:
@@
Utilizing (17), (27), (28) and (28) and combining with the early time result (23), we derive the full-time evolution of the RRE for a primary state and a descendant state
@@
Eq. (30) shows that although generally speaking, the RRE is asymmetrical with respect to the two density matrices that constitute it, RRE indeed can serve as a “measure” of the distinguishability of two states that are indistinguishable in the context of entanglement entropy.
3.3 $k^{\rm th}$ relative Rényi entropy for two general descendant states
We next consider the RRE for two general descendant states, which are generated from acting two different descendant operators belonging to the same conformal family on the vacuum. Let the two density matrices $\rho$ and $\sigma$ take the form of
@@ 3.3 $k^{\rm th}$ relative Rényi entropy for two general descendant states
where $L_{-\{K_{i}\}}\equiv L_{-k_{i_{1}}}L_{-k_{i_{2}}}...L_{-k_{i_{n_{i}}}}$, $(0\leq k_{i_{1}}\leq k_{i_{2}}\leq...\leq k_{i_{n_{i}}})$, and $L_{-\{\bar{K}_{i}\}}\equiv L_{-\bar{k}_{\bar{i}_{1}}}L_{-\bar{k}_{\bar{i}_{2}}}...L_{-\bar{k}_{\bar{i}_{n_{i}}}}$, $(0\leq\bar{k}_{\bar{i}_{1}}\leq\bar{k}_{\bar{i}_{2}}\leq...\leq\bar{k}_{\bar{i}_{n_{i}}})$. Likewise for $L_{-\{K^{\prime}_{j}\}}$ and $L_{-\{\bar{K}^{\prime}_{j}\}}$.
The notation of the form $\text{tr}(\rho\sigma^{k-1})$ now in this case is
$\text{tr}(\rho\sigma^{k-1})=\log$
$\frac{\langle L_{-\{K_{i}\}}\bar{L}_{-\{\bar{K}_{i}\}}\mathcal{O}(w_{1},\bar{w}_{1})L_{-\{K_{i}\}}\bar{L}_{-\{\bar{K}_{i}\}}\mathcal{O}(w_{2},\bar{w}_{2})L_{-\{K^{\prime}_{j}\}}\bar{L}_{-\{\bar{K}^{\prime}_{j}\}}\mathcal{O}(w_{3},\bar{w}_{3})\dots L_{-\{K^{\prime}_{j}\}}\bar{L}_{-\{\bar{K}^{\prime}_{j}\}}\mathcal{O}(w_{2k},\bar{w}_{2k})\rangle_{\Sigma_{k}}}{\langle L_{-\{K_{i}\}}\bar{L}_{-\{\bar{K}_{i}\}}\mathcal{O}(w_{1},\bar{w}_{1})L_{-\{K_{i}\}}\bar{L}_{-\{\bar{K}_{i}\}}\mathcal{O}(w_{2},\bar{w}_{2})\rangle_{\Sigma_{1}}\langle L_{-\{K^{\prime}_{j}\}}\bar{L}_{-\{\bar{K}^{\prime}_{j}\}}\mathcal{O}(w_{1},\bar{w}_{1})L_{-\{K^{\prime}_{j}\}}\bar{L}_{-\{\bar{K}^{\prime}_{j}\}}\mathcal{O}(w_{2},\bar{w}_{2})\rangle_{\Sigma_{1}}^{k-1}}$ (32)
@@
with $|K_{i}|\equiv\sum_{j=1}^{n_{i}}k_{ij}~{}\text{and}~{}|\bar{K}_{i}|\equiv\sum_{j=1}^{\bar{n}_{i}}\bar{k}_{ij}$. The coefficient of the two-point correlation function for generic descendant operators can be evaluated using the algorithm in appendix A and the program in appendix B, and we keep the notation $c_{0}$ and $\bar{c}_{0}$ for convenient in the following content.
@@
Eq.(40) shows that RRE can serve as a “measure” of the distinguishability of two descendant states. Interestingly, even though the states forming the two density matrices $\rho$ and $\sigma$ include both holomorphic and antiholomorphic parts, their RRE ultimately reflects only the information of the holomorphic part. We will discuss the underlying physical picture of this in detail in subsection 4.1. Additionally, note that in RCFTs, the result of the RRE ultimately depends only on a series of coefficients of two-point functions, which allows us to explore the relationship between RRE and some known metrics in the following subsection.
3.4 $k^{\rm th}$ relative Rényi entropy and trace square distance
Although we propose that RRE can serve as a measure to distinguish between different states, it is not a measure since it is not symmetrical. However, from the results above, such as (40), there exist many cases whose RRE are symmetric, arising from the fact that the Rényi entropy $\text{tr}\rho^{k}$ of many states belonging to the same conformal family are universal [28], and the term of the form $\text{tr}\rho\log\sigma$ only depends on the several coefficients of two-point functions (38) where each one is asymmetrical individually, but when combined, they become symmetrical. Therefore, in this subsection, we will explore the relationship between RRE and the trace square distance [57] under the premise of symmetry.
where the factor $\rho_{vacuum}^{2}$, in particular, removes any UV divergences and allows to direct express the TSD in terms of four-point functions on the two-sheeted surface $\Sigma_{2}$ and two-point functions on $\Sigma_{1}$,
Eq. (43) shows that by calculating the trace square distance (TSD), we can obtain the $k^{\rm th}$ RRE. This holds for many cases, as discussed in subsection 3.3, and requires only the computation of correlation functions on a 2-sheeted Riemann surface. However, the examples we have considered thus far do not cover all possibilities. The conditions the reduced density matrices must satisfy for RRE to exhibit symmetry and the relationship between RRE and other traditional metrics remain open questions.
4 Relative Rényi entropy for linear combination operators
where
The $k^{\rm th}$ RRE of $V_{\alpha}$ and $V_{\beta}$ is
which can be evaluated as
$=\frac{1}{k-1}(\log 1-\log 1)=0$ (47)
at the early time according to (11), (33) and (34) while its late time behavior is more complicated. Eq. (24) imply that for $t>\left|x\right|$, the dominant channel for the holomorphic sector of the $2k$-point function on $\Sigma_{1}$ is
Therefore, at the late time, the $k^{\rm th}$ relative Rényi entropy of $V_{\alpha}$ and $V_{\beta}$ (46) can be evaluated as
$S_{k}(\rho_{A}||\sigma_{A})=$
Eq. (50) can be simplified to
As a specific example, we would calculate the RRE for $\partial\mathcal{O}+\bar{\partial}\mathcal{O}$ and ¯$\bar{\partial}\mathcal{O}$ in detail. The coefficient for the two-point function of the primary operator $\mathcal{O}$ has been normalized to $1$, i.e.
Replacing all the matrices in (52) with (54), the RRE for $\partial\mathcal{O}+\bar{\partial}\mathcal{O}$ and ¯$\bar{\partial}\mathcal{O}$ is
4.1 Relative Rényi entropy and quasi-particle
Using the results in subsection 3.3, we can derive the full-time evolution of RRE between some simple reduced density matrices, which are locally excited by $\partial\mathcal{O}$, $\mathcal{O}$, ¯$\bar{\partial}\mathcal{O}$ and ¯$\bar{\partial}^{2}\mathcal{O}$ separately, i.e.
and their results are
We continue to choose the subsystem as $[0,+\infty)$, with the excitation point initially located on the left side of the subsystem. Eqs. (57) and (58) reveal that although the operators forming the density matrices $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{3}$ differ, their reduced density matrices evolve identically over time. From the perspective of relative entropy, a relative entropy of zero implies that the two reduced density matrices are identical. However, does this imply they are indeed identical? Referring to (56), it seems unconvincing to conclude that these density matrices are indeed the same. One possible explanation is that relative entropy may not be reliable for determining whether two reduced-density matrices are distinct.
Unlike in Eq. (58), changing the excitation position of the operator now makes the reduced density matrices, which were previously indistinguishable by relative entropy, distinguishable. Therefore, relative entropy can still be a tool to determine whether two reduced-density matrices are identical. However, this criterion does not ensure the detectability of all density matrices under any excitation, as it depends on the relative position between the excitation point and the subsystem.
For simplicity, let’s set
$\rho^{2}=\frac{\partial\mathcal{O}+\bar{\partial}\mathcal{O}|\Omega\rangle\langle\Omega|\partial\mathcal{O}+\bar{\partial}\mathcal{O}}{\langle(\partial\mathcal{O}+\bar{\partial}\mathcal{O})(\partial\mathcal{O}+\bar{\partial}\mathcal{O})\rangle},\ \mathrm{and}\ \rho^{3}=\frac{1}{2}\frac{\partial\mathcal{O}|\Omega\rangle\langle\Omega|\partial\mathcal{O}}{\langle\partial\mathcal{O}\partial\mathcal{O}\rangle}+\frac{1}{2}\frac{\bar{\partial}\mathcal{O}|\Omega\rangle\langle\Omega|\bar{\partial}\mathcal{O}}{\langle\bar{\partial}\mathcal{O}\bar{\partial}\mathcal{O}\rangle},$ (60)
and it is easy to derive the following relation
where $S_{k}(\rho^{3}_{A})$ represents the $k^{\rm th}$ Rényi entropy of $\rho^{3}_{A}$. Eq. (61) shows that, regardless of whether the reference state is a linear combination state or a mixed state, $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{3}$ remain indistinguishable. The quasi-particle picture explains why relative entropy may sometimes fail to distinguish between density matrices that appear entirely different.
since $\sigma^{1}_{A}$, $\sigma^{2}_{A}$, and $\sigma^{3}_{A}$ contain different anti-holomorphic information. However, operators like $\partial\mathcal{O}$ and $\partial^{2}\mathcal{O}$ become indistinguishable in this case.
5 Relative Rényi entropy and entanglement wedge
The density matrix considered in this section is constructed by a locally excited primary operator $\mathcal{O}(w,\bar{w})$ in a 2D CFT on a complex plane $R^{2}$ where we set $(w,\bar{w})=(x+i\tau,x-i\tau)$, i.e.
$ds^{2}=\eta^{-2}(d\eta^{2}+dx^{2}+d\tau^{2}).$ (65)
Relative entropy is an effective quantum information measure that distinguishes two density matrices. We choose the subsystem $A$ to be an interval $0\leq x\leq L$ at $\tau=0$ for convenient, and the boundary of $M_{A}$ in the CFT is given by
$|w-L/2|=L/2.$ (66)
In this section, we use the sandwiched RRE
$\rho=\frac{\mathcal{O}(w,\bar{w})|\Omega\rangle\langle\Omega|\mathcal{O}^{\dagger}(\bar{w},w)}{\langle\mathcal{O}^{\dagger}(\bar{w},w)\mathcal{O}(w,\bar{w})\rangle},\quad\rho^{\prime}=\frac{\mathcal{O}(w^{\prime},\bar{w}^{\prime})|\Omega\rangle\langle\Omega|\mathcal{O}^{\dagger}(\bar{w}^{\prime},w^{\prime})}{\langle\mathcal{O}^{\dagger}(\bar{w}^{\prime},w^{\prime})\mathcal{O}(w^{\prime},\bar{w}^{\prime})\rangle}.$ (68)
to reproduce the entanglement wedge.
$A_{m,n}=$ $\text{tr}[(\rho^{m}\rho^{\prime}\rho^{m})^{n}]$
and
[…]
and thus
When $\alpha=1/2$, (73) reproduces the fidelity in [53], however, when $\alpha\to 1$, $S_{\alpha}(\rho\mid\mid\rho^{\prime})$ is diverge [^7] When calculating relative entropy, divergence is very common. We will discuss this in section 6. Surprisingly, if $\alpha=2$, (73) would give rise to a finite result, the collision relative entropy [36]
We can expand $w^{\prime}$ and $\bar{w}^{\prime}$ around $w$ and $\bar{w}$ as:
$w^{\prime}=w+\delta dw\quad\bar{w}^{\prime}=\bar{w}+\delta\bar{d}w,$ (75)
$ds^{2}=\frac{2h}{\tau^{2}}(dx^{2}+d\tau^{2}).$ (76)
which is proportional to the Bures metric. Similarly, in a 2D holographic CFT with a circle compactification $x\sim x+2\pi$, we can obtain the metric
$ds^{2}=\frac{2h}{\sinh^{2}\tau}(dx^{2}+d\tau^{2})$ (77)
if $w$ inside the wedge.
Therefore, the collision relative entropy can produce a Bures-like metric proportional to the AdS metric on a time slice and precisely reproduce the expected entanglement wedge from quantum information in 2D holographic CFTs.
6 Conclusions and discussions
This paper investigates the RRE in local operator quenches of RCFTs and holographic CFTs. In RCFTs, the RRE between vacuum and primary operators diverges, as shown in (15), due to limitations of the regularization scheme typically used for entanglement entropy in the vacuum state. For descendant operators within the same conformal family (including primaries as special cases), the RRE often exhibits monotonic time evolution, as illustrated in (40). This behavior can be interpreted by quasi-particles propagating at the speed of light into the subsystem: the time the RRE reaches zero depends on the distance between the initial excitation point and the subsystem, and the RRE’s magnitude depends on the ordering of the excitation operators. In certain instances, the RRE also displays symmetry under specific operator excitations, prompting an investigation into its relationship with the trace squared distance metric (43).
For two linear combination operators, the RRE during time evolution depends solely on certain finite-dimensional matrices (51), whose dimensions correspond to the number of descendant operators in each combination and whose elements depend on combination coefficients and operator orders. Some examples show that despite structural differences in the operators forming distinct density matrices, their RRE can still be zero. The quasi-particle propagation model suggests that relative (Rényi) entropy has limitations in distinguishing reduced density matrices, strongly depending on the position of operator excitations relative to the subsystem. Thus, rather than asserting that relative entropy fails to differentiate between operators, it is more accurate to state that it only distinguishes information from operators that have entered or exited the subsystem.
Finally, we reconstruct the geometry of the entanglement wedge from a quantum information perspective. In holographic CFTs, collision relative entropy induces a Bruś-like metric, ultimately revealing the sharp structure discussed in [53], which provides the geometric structure of the entanglement wedge.
It is important to note that the RRE studied in this paper cannot be analytically continued in either RCFTs or holographic CFTs. Analytical continuation remains challenging, and the reasons for this limitation are not yet fully understood. We hope to gain deeper insights into this problem in future work. Additionally, since the RRE is sometimes symmetric between two density matrices, exploring the sufficient and necessary conditions that lead to this symmetry remains an interesting open question. We would like to leave further detailed investigation on them to future work.
Acknowledgements
We thank Rohit Mishra, Masahiro Nozaki, Yuan Sun, Long Zhao, and Yang Liu for valuable discussions on this work. H. O., H. Z., and Z. Z. are supported by the Science and Technology Development Plan Project of Jilin Province, China Grant No. 20240101326JC. S. H. acknowledges financial support from the Max Planck Partner Group and the Natural Science Foundation of China Grants No. 12075101, No. 12235016, No. 12347209, and No. 12475053. H. O. is also supported by the National Natural Science Foundation of China Grant No. 12205115.
Appendix A Correlation function of two descendant operators

### v2

@@ Appendix A Useful formulae for computing the k-th RRE
For convenience, unless otherwise specified, we select the subsystem $A$ to be $0,\infty)$ in our paper, and the 2kk points $z_{1},z_{2},...,z_{2k}$ in the $z$-coordinates are given by
@@
Appendix B Relative Rényi entropies for two locally excited states
This section presents the details of the $k^{\text{th}}$ relative Rényi entropy growth calculation for two locally excited states within the same conformal family.
B.1 $|\Omega\rangle$ and $\mathcal{O}(x)|\Omega\rangle$
Two density matrices in this case are given by
@@ Appendix B Relative Rényi entropies for two locally excited states
The second term on the RHS of (62), by utilizing the conformal map (58), can be reformulated as
$\text{tr}(\rho\sigma^{k-1})=\frac{\left|\frac{dw_{1}}{dz_{1}}\right|^{-2\Delta}\left|\frac{dw_{2}}{dz_{2}}\right|^{-2\Delta}\langle\mathcal{O}(z_{1},\bar{z}_{1})\mathcal{O}(z_{2},\bar{z}_{2})\mathbbm{1}\dots\mathbbm{1}\rangle_{\Sigma_{1}}}{\langle\mathcal{O}(w_{1},\bar{w}_{1})\mathcal{O}(w_{2},\bar{w}_{1})\rangle_{\Sigma_{1}}},$ (63)
At early times, we had
@@
$\text{tr}(\rho\sigma^{k-1})=\frac{k^{-4\Delta}(x^{2}-t^{2})^{\frac{-2(k-1)\Delta}{k}}k^{4\Delta}(x^{2}-t^{2})^{\frac{2(k-1)\Delta}{k}}\left|2\epsilon\right|^{-4\Delta}}{\left|2\epsilon\right|^{-4\Delta}}=0.$ (65)
So, the RRE of the primary and identity operators was zero at early times.
Next, we would discuss the late-time behavior of (63). The numerator of the logarithm function can be evaluated as
@@
Therefore, at late times, the $k^{\rm th}$ RRE of primary operator and identity operator saturates to
@@
B.2 $L_{-n}\mathcal{O}(x)|\Omega\rangle$ and $\mathcal{O}(x)|\Omega\rangle$
The two density matrices are constructed from a primary operator and a descendant operator belonging to the same conformal family, i.e.
@@
From (70) and (71) and replacing all the coordinates with (60), at early times, we find the two-point functions on $\Sigma_{k}$ and on $\Sigma_{1}$ are the same, i.e.
@@
Therefore, combining (64), (14), (70) and (74), the $k^{\rm th}$ RRE of the primary operator and descendant operator at early times is
@@
Based on (60), it can be found that when $t>\left|x\right|$, the $2k$ holomorphic coordinates and the $2k$ anti-holomorphic coordinates approach each other in distinct pairings [106]
@@
$=$ $\langle\mathcal{O}(z_{1})L_{-n}\mathcal{O}(z_{4})\dots L_{-n}\mathcal{O}(z_{2j+1})L_{-n}\mathcal{O}(z_{2j+4})\dots L_{-n}\mathcal{O}(z_{2k-1})\mathcal{O}(z_{2})\rangle_{\Sigma_{1}}$
$\times$ $\langle\mathcal{O}(\bar{z}_{1})\mathcal{O}(\bar{z}_{2})\mathcal{O}(\bar{z}_{3})\dots\mathcal{O}(\bar{z}_{2k})\rangle_{\Sigma_{1}}$
@@
$\times\langle\mathcal{O}(\bar{z}_{1})\mathcal{O}(\bar{z}_{2})\rangle_{\Sigma_{1}}\dots\langle\mathcal{O}(\bar{z}_{2k-3})\mathcal{O}(\bar{z}_{2k-2})\rangle_{\Sigma_{1}}\langle\mathcal{O}(\bar{z}_{2k-1})\mathcal{O}(\bar{z}_{2k})\rangle_{\Sigma_{1}}$ (77)
@@
As pointed out in [105], at the late time, the holomorphic part of the two-point function for two operators belonging to the same conformal family on $\Sigma_{k}$ and $\Sigma_{1}$ has the following relation:
@@
Utilizing (14), (79), (80), and (80) and combining with the early time result (75), we derive the full-time evolution of the RRE for a primary state and a descendant state
@@
B.3 Two general descendant states
Let the two density matrices $\rho$ and $\sigma$ take the form of
@@ B.3 Two general descendant states
with $|K_{i}|\equiv\sum_{j=1}^{n_{i}}k_{ij}~\text{and}~|\bar{K}_{i}|\equiv\sum_{j=1}^{\bar{n}_{i}}\bar{k}_{ij}$. The coefficient of the two-point correlation function for generic descendant operators can be evaluated using the algorithm in Appendix C and the program in Appendix D. We retain the notation $c_{0}$ and $\bar{c}_{0}$ for convenience in the following content.
@@
Appendix C Correlation function of two descendant operators